# Learning to Mask and Permute Visual Tokens for Vision Transformer Pre-Training

### v1

@@ 4.2 $k$-CLIP: Discretized CLIP-based Tokenizer
The role of visual tokenizers in pre-training pipelines is significant, as they provide crucial guidance for downstream fine-tuning outcomes. To possibly reduce the overhead of tokenizer retraining of previously proposed approaches peng2022beit , we explore the impact of directly utilizing discretized CLIP features. In particular, we propose a novel visual tokenizer, called $k$-CLIP, that employs discretized CLIP features as visual tokens. Our method does not rely on any pre-training or supervised data to create the visual tokens, thus enabling pre-training without access to large amounts of labeled data or a particular pre-training dataset. Furthermore, the use of CLIP features enables our method to capture high-level visual semantics that are more meaningful for downstream tasks, further improving the performance of the learned representations.
Specifically, we sample visual features from the ImageNet dataset deng2009imagenet using the CLIP model radford2021learning and cluster them using $k$-means to obtain $|\mathcal{V}|=8192$ centroids. During pre-training, these centroids are indexed to retrieve the corresponding visual tokens for prediction. Formally, our visual tokenizer $\mathcal{T}(\bm{x})$ consists of a ResNet-based CLIP visual encoder and a $k$-means model. The visual encoder $f_{v}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{H_{c}\times W_{c}\times D_{c}}$ maps an input image $\bm{x}$ to a grid of visual features that correspond to the activations produced by the last convolutional layer of the CLIP backbone. The visual features are then reshaped to a $N_{c}\times D_{c}$ matrix, where $N_{c}=H_{c}\times W_{c}$. In our method, the CLIP visual encoder generates visual features of shape ${N_{c}\times D_{c}}={196\times 4096}$, which are subsequently indexed by $k$-means and mapped to a sequence of $196$ discrete visual tokens $\bm{v}=[v_{1},\dots,v_{N}]\in\mathcal{I}$. The set of visual tokens $\bm{v}=\{v_{i}\}_{i=1}^{N}$ is defined by the $k$-means centroid indexes =$\mathcal{I}=\{1,\dots,|\mathcal{V}|\}$. To mitigate the computational complexity, we randomly sample $2\%$ of the approximately $250$ million $D_{c}$-dimensional CLIP features extracted from ImageNet. The sampled feature collection is then used to fit the $k$-means clustering model.
@@ 5.1 Experimental Setup
Self-supervised pre-training literature has introduced several training and fine-tuning procedures that encompass different tokenizers, hyperparameters, and visual backbones. To ensure a fair and unbiased comparison among these pre-training algorithms, we opt to evaluate our MaPeT method, as well as the related approaches in the literature, using identical experimental configurations. This strategy allows for the isolation of the algorithmic factor in the experiments and promotes unbiased comparisons of pre-training objectives. In order to assess the effectiveness of the pre-trained objective under consideration, we fine-tune our models on a downstream classification task.
Pre-training setup. We first investigate the influence of the visual tokenizer, evaluating the proposed $k$-CLIP against VQ-KD peng2022beit . To minimize the computational effort, we pre-process the ImageNet-1k training dataset deng2009imagenet for both the tokenizers and store the visual tokens associated with each image. For this reason, our pre-training augmentation policy only includes color jittering to preserve image patch positions corresponding to the pre-extracted visual tokens.
We compare MaPeT against different pre-training objectives. In particular, we employ a ViT-based backbone that is pre-trained according to the pure MIM objective formulation. Note that this differs from the pre-training strategy proposed by the BEiT approach bao2022beit as it lacks the block-wise masking algorithm, which progressively extracts multiple blocks of patches until 40% of the positions are masked. In our MIM-based pre-training, we replace the block-wise masking strategy with a random patch masking approach, thus keeping it similar to a BERT-like solution devlin2019bert applied to Computer Vision tasks. Analogously, we also pre-train a ViT backbone through a double-stream architecture according to the PIM objective described in Sec. 3. Moreover, we consider the standard BEiT model bao2022beit (i.e., a MIM-based pre-training with block-wise masking) and CAE chen2022context . The CAE method employs an encoder-decoder architecture where the encoder processes only visible image patches (50% of the entire image), while the remaining 50% is masked. A latent contextual regressor predicts the masked representation based on the encoder output, and a lightweight decoder processes the output of the regressor, which is then used to predict the visual token of the related masked patches.
@@
Table 2: Fine-tuning results in comparison with existing self-supervised pre-training approaches, in terms of top-1 and linear probe accuracy on ImageNet-1k.
@@
ViT dosovitskiy2021image - 73.7 - 79.8 - 81.8 -
CAE chen2022context VQ-KD 73.4 52.4 81.6 63.4 83.5 69.4
BEiT bao2022beit VQ-KD 75.0 62.0 82.2 72.6 84.4 75.0
@@
CAE chen2022context $k$-CLIP 73.7 53.9 81.5 62.9 82.7 67.6
BEiT bao2022beit $k$-CLIP 74.8 61.5 81.9 71.1 83.3 73.3
@@
Image classification setup. During the classification fine-tuning stage, the final hidden layer of the ViT-based backbone extracts features that are then combined via average pooling to generate a global image representation. This representation is subsequently fed into a softmax classifier. Following the linear probing experiment reported in bao2022beit , we also train a linear classifier head over the output representation produced by the frozen pre-trained backbone.
We design three different model variants based on Vision Transformer dosovitskiy2021image , i.e., ViT-Tiny (ViT-T), ViT-Small (ViT-S), and ViT-Base (ViT-B). Our MaPeT model is trained by setting the cutting point $c$ to $50$, $50$, and $60$, respectively. We refer the reader to the supplementary material for detailed pre-training and fine-tuning hyperparameters, and for an ablation analysis on cutting point values.
@@ 5.2 Experimental Results
BEiT bao2022beit 300 81.7 82.9
CAE chen2022context 300 82.0 83.6
SplitMask el2021large 300 - 83.6
MaskFeat wei2022masked 300 - 83.6
PeCo dong2023peco 300 - 84.1
MVP wei2022mvp 300 - 84.4
BEiT v2 peng2022beit 300 - 85.0
@@
ViT-B Method # Epochs Top-1 Acc. (%) BEiT bao2022beit 800 83.2 CAE chen2022context 800 83.8 CAE chen2022context 1600 83.9 BEiT v2 peng2022beit 1600 85.5
@@
We also report the comparison of our best variants and other state-of-the-art self-supervised pre-training models in Table 3. Note that the variability in the experimental settings and the different supervisory signals used in each approach may affect the fairness of the comparisons. Notably, although current literature has not extensively explored performance benchmarking on the ViT-S backbone, the Small version of our MaPeT model outperforms BEiT bao2022beit and CAE chen2022context by $0.5\%$ and $0.2\%$ respectively, yielding the best performance for both $k$-CLIP and VQ-KD. In the case of the ViT-B backbone, MaPeT surpasses the majority of the considered approaches especially when using the VQ-KD tokenizer, except for BEiT v2 peng2022beit which gets slightly better results. This performance gap can be explained by the lack of data augmentation in our model pre-training, which can significantly increase performance but at a higher computational cost.
@@
All the results presented in Table 3 are obtained after pre-training the models for 300 epochs. For completeness, on the right side, we report the performance of other methods when pre-trained for a considerably large number of epochs (i.e., 800 and 1600). While a direct comparison using different pre-training epochs may not be completely informative, MaPeT still proves to perform better than other methods pre-trained for a larger number of epochs such as BEiT and CAE.
5.3 Visual Tokenizer Analysis
In this section, we present a comprehensive qualitative and quantitative analysis of our proposed $k$-CLIP and compare its performance to the existing DALL-E ramesh2021zero and VQ-KD peng2022beit tokenizers.
Image classification with discrete visual tokens. In this section, we analyze image classification performance when using discretized visual tokens directly as input to the model. We conduct this analysis using the previously employed tokenizers (i.e., our $k$-CLIP and VQ-KD peng2022beit ) and also include the DALL-E visual tokenizer ramesh2021zero , which consists of a discrete variational autoencoder model. Specifically, we map the sequence of visual tokens to an embedding space of dimension $D$ through learnable embeddings trained in conjunction with two different model backbones: a lightweight MLP classification head and a ViT-Tiny model. The MLP head comprises an embedding layer, a linear layer with output dimension $D=192$, a ReLU activation, an average pooling operation, and an additional linear layer that projects the pooled features to the number of classes (i.e., 1000).
Figure 3: Visualization of image patches corresponding to the discrete tokens of our $k$-CLIP codebook, where semantically similar patches consistently share the same discrete token.
Table 4: Accuracy of image classification when employing visual tokens as model input. Note that using CLIP features for zero-shot image classification leads to a top-1 accuracy of 73.6%.
MLP ViT-T
Tokenizer Top-1 Acc. (%) Top-5 Acc. (%) Top-1 Acc. (%) Top-5 Acc. (%)
DALL-E 4.1 11.7 9.1 21.9
VQ-KD 68.1 89.8 72.3 92.2
$\mathbf{k}$-CLIP 72.8 93.2 76.2 94.9
6 Conclusion
This paper presents MaPeT, a novel self-supervised pre-training approach designed for vision tasks. Our model effectively tackles the limitations of standard Masked Image Modeling, employing a permutation-based objective to capture the interdependencies among predicted tokens and auxiliary position information to enable the model to access a full sequence of image patches. Moreover, we introduce the $k$-CLIP tokenizer that can densely capture the semantic information of the visual input by leveraging discretized CLIP features as visual tokens. Experimental results demonstrate that our approach achieves improved fine-tuning performance on image classification, outperforming direct competitors under the same experimental setting.
Acknowledgments and Disclosure of Funding
We thank CINECA for providing computational resources. This work has partially been supported by the projects PNRR-M4C2 (PE00000013) “FAIR - Future Artificial Intelligence Research” funded by the European Commission and “ELSA - European Lighthouse on Secure and Safe AI” funded by the EU (GA 101070617).
Supplementary Material
In the following, we explore the adaptability of our model to novel data domains and examine the influence of varying the reconstruction factor on the performance of the model. Furthermore, we provide additional qualitative visualizations of the codebook associated with the proposed $k$-CLIP tokenizer, supplementary implementation details, and a discussion on the potential limitations and social implications of the proposed methodology.
A Cross-Domain Transfer Learning Capabilities
In this section, we examine the generalization capabilities of our proposed self-supervised pre-training technique, in comparison to BEiT bao2022beit and CAE chen2022context . All the considered architectures employ the VQ-KD visual tokenizer peng2022beit and undergo pre-training on the extensive ImageNet-1k dataset deng2009imagenet . Subsequently, they are fine-tuned on three distinct data domains, namely Stanford-Cars krause20133d , Food-101 bossard2014food , and FGVC-Aircraft maji13fine-grained . These datasets are chosen to represent diverse and real-world scenarios, ranging from object recognition in the automotive domain to food and aircraft classification. By employing a linear probe evaluation, we can quantitatively measure the ability of our pre-trained model to transfer knowledge and adapt to new tasks without fine-tuning.
The results, presented in Table 5, clearly demonstrate the superior performance of our MaPeT model across all considered datasets. Our model outperforms both BEiT and CAE, highlighting its robustness and efficacy in capturing meaningful visual representations.
These findings not only underline the potential of MaPeT as a powerful pre-training technique but also emphasize its cross-domain transfer learning capabilities, which enable practical relevance in various real-world applications.
B Reconstruction Ratio Analysis
Here we discuss the relationship between the reconstruction ratio employed in the MaPeT model and its impact on image classification performance. The reconstruction ratio indicates the proportion of target patches in relation to the entire input sequence.
An advantage of MaPeT is that it attends to the target patch $\bm{x_{z_{t}}}\in\bm{x}_{z_{>c}}$ during the prediction of $\bm{x_{z_{t+1}}}$, allowing for the cutting point $c$ to be positioned at every value in the interval $c\in[1\ldots N-1]$. This unique feature of MaPeT enables the potential reconstruction of all patches in an incremental and randomized manner, from the first to the last patch. On the other hand, the BEiT and CAE models lack the ability to reconstruct all image patches because doing so would lead to a masking ratio that is too high, resulting in the loss of visual context necessary for the reconstruction of the masked elements. Table 6 refers to the analysis of the reconstruction ratio on the ViT-B architecture. It can be noted that MaPeT exhibits a preference for a reconstruction ratio of approximately 70% of the complete sequence. Indeed, the initial target patches are compelled to establish correlations with a small portion of visible patches (namely, 30% of the entire sequence) randomly distributed throughout the image.
The model can acquire long-range spatial dependencies that are contingent on the position of the target patch in the permuted order. As the final patches attend to the majority of the visual content, they are likely to concentrate more on neighboring image patches, consistent with the principle of spatial locality. A reduction in the reconstruction ratio leads to an increase in the number of visible patches attended by the target patches that occur early in the permuted order. This results in an increase in the likelihood of having neighboring visible patches, which, in turn, diminishes the learned long-range spatial dependencies. In contrast, an excessively high reconstruction ratio of 85% makes the pre-training objective of MaPeT excessively difficult, as early predictions are likely to be arbitrary, having access to only 15% of the overall visual information. Consequently, excessively high or low reconstruction ratios impair the performance, as evidenced by Table 6.
Table 5: Comparison of model performance on cross-domain transfer learning. We measure the linear probe accuracy of our proposed MaPeT model compared to BEiT and CAE on three distinct data domains: Stanford-Cars, Food-101, and FGVC-Aircraft.
@@ B Reconstruction Ratio Analysis
Method Tokenizer Stanford-Cars krause20133d Food-101 bossard2014food FGVC-Aircraft maji13fine-grained
CAE chen2022context VQ-KD 53.6 81.4 40.5
BEiT bao2022beit VQ-KD 64.5 86.9 44.9
@@
Table 6: Ablation study on the impact of the reconstruction factor (%) on the top-1 fine-tuning accuracy of pre-trained ViT-B under the VQ-KD setting.
Cutting point $c$ Reconstruction ratio Tokenizer Top-1 Accuracy (%)
@@
These observations shed light on the interplay between the reconstruction ratio and the capacity of the model to capture spatial relationships. The optimal reconstruction ratio, as evidenced by our findings, strikes a balance between capturing long-range dependencies and leveraging the spatial locality inherent in the visual data.
C Additional Codebook Visualization

### v2

@@ 4.2 $k$-CLIP: Discretized CLIP-based Tokenizer
The role of visual tokenizers in pre-training pipelines is significant, as they provide crucial guidance for downstream fine-tuning outcomes. To possibly reduce the overhead of tokenizer retraining of previously proposed approaches (Peng et al., 2022), we explore the impact of directly utilizing discretized CLIP features. In particular, we propose a novel visual tokenizer, called $k$-CLIP, that employs discretized CLIP features as visual tokens. Our method does not rely on any pre-training or supervised data to create the visual tokens, thus enabling pre-training without access to large amounts of labeled data or a particular pre-training dataset. Furthermore, the use of CLIP features enables our method to capture high-level visual semantics that are more meaningful for downstream tasks, further improving the performance of the learned representations.
Specifically, we sample visual features from the ImageNet dataset (Deng et al., 2009) using the CLIP model (Radford et al., 2021) and cluster them using $k$-means to obtain $|\mathcal{V}|=8192$ centroids. During pre-training, these centroids are indexed to retrieve the corresponding visual tokens for prediction. Formally, our visual tokenizer $\mathcal{T}(\bm{x})$ consists of a ResNet-based CLIP visual encoder and a $k$-means model. The visual encoder $f_{v}:\mathbb{R}^{H\times W\times C}\rightarrow\mathbb{R}^{H_{c}\times W_{c}\times D_{c}}$ maps an input image $\bm{x}$ to a grid of visual features that correspond to the activations produced by the last convolutional layer of the CLIP backbone. The visual features are then reshaped to a $N_{c}\times D_{c}$ matrix, where $N_{c}=H_{c}\times W_{c}$. In our method, the CLIP visual encoder generates visual features of shape ${N_{c}\times D_{c}}={196\times 4096}$, which are subsequently indexed by $k$-means and mapped to a sequence of $196$ discrete visual tokens $\bm{v}=[v_{1},\dots,v_{N}]\in\mathcal{I}$. The set of visual tokens $\bm{v}=\{v_{i}\}_{i=1}^{N}$ is defined by the $k$-means centroid indexes =$\mathcal{I}=\{1,\dots,|\mathcal{V}|\}$. To mitigate the computational complexity, we randomly sample $2\%$ of the approximately $250$ million $D_{c}$-dimensional CLIP features extracted from ImageNet. The sampled feature collection is then used to fit the $k$-means clustering model.
@@ 5.1 Experimental Setup
Self-supervised pre-training literature has introduced several training and fine-tuning procedures that encompass different tokenizers, hyperparameters, and visual backbones. To ensure a fair and unbiased comparison among these pre-training algorithms, we opt to evaluate our MaPeT method, as well as the related approaches in the literature, using identical experimental configurations. This strategy allows for the isolation of the algorithmic factor in the experiments and promotes unbiased comparisons of pre-training objectives. To assess the effectiveness of the pre-trained objective under consideration, we fine-tune our models on a downstream classification task. Additionally, following recent literature, we perform fine-tuning experiments for semantic segmentation, to assess the generalization of the proposed pre-training objective to dense predictions.
@@
ViT (Dosovitskiy et al., 2021) - 73.7 - 79.8 - 81.8 -
CAE (Chen et al., 2023) VQ-KD 73.4 52.4 81.6 63.4 83.5 69.4
BEiT (Bao et al., 2022) VQ-KD 75.0 62.0 82.2 72.6 84.4 75.0
@@
CAE (Chen et al., 2023) $k$-CLIP 73.7 53.9 81.5 62.9 82.7 67.6
BEiT (Bao et al., 2022) $k$-CLIP 74.8 61.5 81.9 71.1 83.3 73.3
@@
Table 2: Fine-tuning results compared to other self-supervised pre-training approaches, in terms of top-1 and linear probe accuracy.
Pre-training setup. We first investigate the influence of the visual tokenizer and evaluate the proposed $k$-CLIP against VQ-KD (Peng et al., 2022). To minimize the computational effort, we pre-process the ImageNet-1k training dataset for both the tokenizers and store the visual tokens associated with each image. For this reason, our pre-training augmentation policy only includes color jittering to preserve image patch positions corresponding to the pre-extracted visual tokens.
We compare MaPeT against different pre-training objectives. In particular, we employ a ViT-based backbone that is pre-trained according to the pure MIM objective formulation. Note that this differs from the pre-training strategy proposed by the BEiT approach (Bao et al., 2022) as it lacks the block-wise masking algorithm, which progressively extracts multiple blocks of patches until 40% of the positions are masked. In our MIM-based pre-training, we replace the block-wise masking strategy with a random patch masking approach, thus keeping it similar to a BERT-like solution (Devlin et al., 2019) applied to Computer Vision tasks. Analogously, we also pre-train a ViT backbone through a double-stream architecture according to the PIM objective described in Sec. 3. Moreover, we consider the standard BEiT model (Bao et al., 2022) (i.e., a MIM-based pre-training with block-wise masking) and CAE (Chen et al., 2023). The CAE method employs an encoder-decoder architecture where the encoder processes only visible image patches (50% of the entire image), while the remaining 50% is masked. A latent contextual regressor predicts the masked representation based on the encoder output, and a lightweight decoder processes the output of the regressor, which is then used to predict the visual token of the related masked patches.
Image classification setup. During the classification fine-tuning stage, the final hidden layer of the ViT-based backbone extracts features that are then combined via average pooling to generate a global image representation. This representation is subsequently fed into a softmax classifier. Following the linear probing experiment reported in (Bao et al., 2022), we also train a linear classifier head over the output representation produced by the frozen pre-trained backbone.
We design three different model variants based on Vision Transformer (Dosovitskiy et al., 2021), i.e., ViT-Tiny (ViT-T), ViT-Small (ViT-S), and ViT-Base (ViT-B). Our MaPeT model is trained by setting the cutting point $c$ to $50$, $50$, and $60$, respectively. We refer the reader to the supplementary material for detailed pre-training and fine-tuning hyperparameters.
Semantic segmentation setup. Semantic segmentation is a pixel-wise classification task that predicts semantic labels for each input image pixel. Our experimental framework follows the setting proposed in BEiT v2 (Peng et al., 2022) and utilizes the ADE20K dataset (Zhou et al., 2017, 2019) as a comprehensive benchmark comprising 25k images spanning 150 semantic categories. For the segmentation architecture, we employ UperNet (Xiao et al., 2018) task layer and fine-tune the model for 160k iterations with input images resized to a resolution of $512\times 512$. The results are reported in terms of mean Intersection over Union (mIoU). To facilitate reproducibility, we include a detailed list of hyperparameters in the supplementary material.
Computational requirements. The pre-training experiments conducted with our MaPeT model involve the utilization of different GPU configurations. Specifically, the ViT-T, ViT-S, and ViT-B models require the deployment of 16, 32, and 64 GPUs, respectively. The pre-training process for each model takes approximately one day to complete. In contrast, during the fine-tuning phase, we employ 4 GPUs for the ViT-T and ViT-S models, and 16 GPUs for the ViT-B model. The fine-tuning duration for each model is 48 hours, 36 hours, and 12 hours, respectively. For all experiments, we utilize an NVIDIA V100 GPU architecture with 16GB of memory.
@@ 5.2 Experimental Results
BEiT (Bao et al., 2022) 300 81.7 82.9
CAE (Chen et al., 2023) 300 82.0 83.6
SplitMask (El-Nouby et al., 2021) 300 - 83.6
MaskFeat (Wei et al., 2022a) 300 - 83.6
PeCo (Dong et al., 2023) 300 - 84.1
MVP (Wei et al., 2022b) 300 - 84.4
BEiT v2 (Peng et al., 2022) 300 - 85.0
@@
BEiT (Bao et al., 2022) 800 - 83.2
CAE (Chen et al., 2023) 800 - 83.8
CAE (Chen et al., 2023) 1600 - 83.9
BEiT v2 (Peng et al., 2022) 1600 - 85.5
@@
We also report the comparison of our best variants and other state-of-the-art self-supervised pre-training models in Table 3. Note that the variability in the experimental settings and the different supervisory signals used in each approach may affect the fairness of the comparisons. Notably, although current literature has not extensively explored performance benchmarking on the ViT-S backbone, the Small version of our MaPeT model outperforms BEiT (Bao et al., 2022) and CAE (Chen et al., 2023) by $0.5\%$ and $0.2\%$ respectively, yielding the best performance for both $k$-CLIP and VQ-KD. In the case of the ViT-B backbone, MaPeT surpasses most of the considered approaches especially when using the VQ-KD tokenizer, except for BEiT v2 (Peng et al., 2022) which gets slightly better results. This performance gap can be explained by the lack of data augmentation in our model pre-training, which can significantly increase performance but at a higher computational cost.
@@
The results in Table 3 are obtained after pre-training the models for 300 epochs. For completeness, in the lower part, we report the performance of other methods when pre-trained for a considerably large number of epochs (i.e., 800 and 1600). While a direct comparison using different pre-training epochs may not be completely informative, MaPeT still proves to perform better than other methods pre-trained for a larger number of epochs such as BEiT and CAE.
mIoU (%)
Method Tokenizer ViT-T ViT-S ViT-B
MIM VQ-KD 39.0 45.6 49.3
PIM VQ-KD 40.2 46.1 49.0
MaPeT VQ-KD 39.3 47.0 50.4
CAE (Chen et al., 2023) VQ-KD 39.2 45.9 50.2
BEiT (Bao et al., 2022) VQ-KD 39.9 46.7 50.3
MaPeT VQ-KD 39.3 47.0 50.4
Table 4: Semantic segmentation results in terms of mIoU on ADE20K, comparing MaPeT with alternative pre-training objectives (top) and other self-supervised pre-training approaches (bottom).
@@
Method Tokenizer Stanford-Cars Food-101 FGVC-Aircraft
CAE (Chen et al., 2023) VQ-KD 53.6 81.4 40.5
BEiT (Bao et al., 2022) VQ-KD 64.5 86.9 44.9
@@
Table 5: Comparison of model performance on cross-domain transfer learning. We measure the linear probe accuracy of our proposed MaPeT model compared to BEiT and CAE on three distinct data domains: Stanford-Cars (Krause et al., 2013), Food-101 (Bossard et al., 2014), and FGVC-Aircraft (Maji et al., 2013).
Semantic segmentation results. To evaluate the effectiveness of MaPeT on dense prediction tasks, we present its performance on the semantic segmentation benchmark ADE20K using the UperNet framework with the VQ-KD tokenizer. Table 4 presents the mIoU scores achieved with various ViT backbones and pre-training strategies. Notably, MaPeT surpasses both MIM and PIM objectives, delivering significant improvements of $+1.4\%$ and $+0.9\%$ on ViT-S, and $+1.1\%$ and $+1.4\%$ on ViT-B, respectively. Compared to CAE (Chen et al., 2023) and BEiT (Bao et al., 2022), MaPeT achieves the highest mIoU across ViT-S and ViT-B backbones, while maintaining competitive performance on ViT-T. These results underscore the superiority of MaPeT over existing pre-training objectives and self-supervised approaches, particularly in addressing positional inconsistencies during pre-training and fine-tuning, which is critical for dense tasks like semantic segmentation. The robust generalization demonstrated by MaPeT highlights its potential as an effective pre-training strategy for segmentation tasks.
Cross-domain transfer learning. As an additional analysis, we examine the generalization capabilities of our proposed self-supervised pre-training technique, compared to BEiT (Bao et al., 2022) and CAE (Chen et al., 2023). All the considered architectures employ the VQ-KD visual tokenizer (Peng et al., 2022) and undergo pre-training on the ImageNet-1k dataset (Deng et al., 2009). Subsequently, they are fine-tuned on three distinct data domains, namely Stanford-Cars (Krause et al., 2013), Food-101 (Bossard et al., 2014), and FGVC-Aircraft (Maji et al., 2013). These datasets are chosen to represent diverse and real-world scenarios, ranging from object recognition in the automotive domain to food and aircraft classification. By employing a linear probe evaluation, we can quantitatively measure the ability of our pre-trained model to transfer knowledge and adapt to new tasks without fine-tuning.
The results, presented in Table 5, clearly demonstrate the superior performance of MaPeT across all considered datasets. Our model outperforms both BEiT and CAE, highlighting its robustness and efficacy in capturing meaningful visual representations. These findings not only underline the potential of MaPeT as a powerful pre-training technique but also emphasize its cross-domain transfer learning capabilities, which enable practical relevance in various real-world applications.
Reconstruction ratio analysis. Here we discuss the relationship between the reconstruction ratio employed in MaPeT and its impact on image classification results. The reconstruction ratio indicates the proportion of target patches in relation to the entire input sequence. An advantage of MaPeT is that it attends to the target patch $\bm{x_{z_{t}}}\in\bm{x}_{z_{>c}}$ during the prediction of $\bm{x_{z_{t+1}}}$, allowing for the cutting point $c$ to be positioned at every value in the interval $c\in[1\ldots N-1]$. This unique feature of MaPeT enables the potential reconstruction of all patches in an incremental and randomized manner, from the first to the last patch. On the other hand, models like BEiT and CAE lack the ability to reconstruct all image patches because this would lead to a masking ratio that is too high, resulting in the loss of visual context necessary for the reconstruction of the masked elements.
Top-1
Cutting Point $c$ Reconstruction Ratio Tokenizer Accuracy (%)
@@
Table 6: Ablation study on the impact of the reconstruction factor on the top-1 fine-tuning accuracy of pre-trained ViT-B under the VQ-KD setting.
Table 6 refers to the analysis of the reconstruction ratio on the ViT-B architecture. It can be noted that MaPeT exhibits a preference for a reconstruction ratio of approximately 70% of the complete sequence. Indeed, the initial target patches are compelled to establish correlations with a small portion of visible patches (namely, 30% of the entire sequence) randomly distributed throughout the image. The model can acquire long-range spatial dependencies that are contingent on the position of the target patch in the permuted order. As the final patches attend to the majority of the visual content, they are likely to concentrate more on neighboring image patches, consistent with the principle of spatial locality. A reduction in the reconstruction ratio leads to an increase in the number of visible patches attended by the target patches that occur early in the permuted order. This results in an increase in the likelihood of having neighboring visible patches, which, in turn, diminishes the learned long-range spatial dependencies. In contrast, an excessively high reconstruction ratio of 85% makes the pre-training objective of MaPeT excessively difficult, as early predictions are likely to be arbitrary, having access to only 15% of the overall visual information. Consequently, excessively high or low reconstruction ratios impair the performance.
Fig. 3: Visualization of image patches corresponding to the discrete tokens of our $k$-CLIP codebook, where semantically similar patches consistently share the same discrete token.
5.3 Visual Tokenizer Analysis
In this section, we present a comprehensive qualitative and quantitative analysis of our proposed $k$-CLIP and compare its performance to the existing DALL-E (Ramesh et al., 2021) and VQ-KD (Peng et al., 2022) tokenizers.
Image classification with discrete visual tokens. In this section, we analyze image classification performance when using discretized visual tokens directly as input to the model. We conduct this analysis using the previously employed tokenizers (i.e., our $k$-CLIP and VQ-KD (Peng et al., 2022)) and also include the DALL-E visual tokenizer (Ramesh et al., 2021), which consists of a discrete variational autoencoder model. Specifically, we map the sequence of visual tokens to an embedding space of dimension $D$ through learnable embeddings trained in conjunction with two different model backbones: a lightweight MLP classification head and a ViT-Tiny model. The MLP head comprises an embedding layer, a linear layer with output dimension $D=192$, a ReLU activation, an average pooling operation, and an additional linear layer that projects the pooled features to the number of classes (i.e., 1000).
MLP ViT-T
Top-1 Top-5 Top-1 Top-5
Tokenizer Acc. (%) Acc. (%) Acc. (%) Acc. (%)
DALL-E 4.1 11.7 9.1 21.9
VQ-KD 68.1 89.8 72.3 92.2
$\mathbf{k}$-CLIP 72.8 93.2 76.2 94.9
Table 7: Accuracy of image classification when employing visual tokens as model input. Note that using CLIP features for zero-shot image classification lead to a top-1 accuracy of 73.6%.
6 Conclusion
This paper presents MaPeT, a novel self-supervised pre-training approach designed for vision tasks. Our model effectively tackles the limitations of standard Masked Image Modeling by employing a permutation-based objective to capture the interdependencies among predicted tokens and auxiliary position information to enable the model to access a full sequence of image patches, thus reducing the discrepancy between pre-training and fine-tuning. Moreover, we introduce the $k$-CLIP tokenizer that can densely capture the semantic information of the visual input by leveraging discretized CLIP features as visual tokens. Experimental results demonstrate that our approach achieves improved fine-tuning performance on image classification and semantic segmentation, outperforming direct competitors under the same experimental setting.
Acknowledgments
We acknowledge the CINECA award under the ISCRA initiative, for the availability of high-performance computing resources. This work has partially been supported by the PNRR-M4C2 (PE00000013) project “FAIR - Future Artificial Intelligence Research” and the Horizon Europe project “ELSA - European Lighthouse on Secure and Safe AI” (GA 101070617), both funded by the European Union.
Supplementary Material
In the following, we offer qualitative visualizations of the codebook associated with the proposed $k$-CLIP tokenizer. Furthermore, we provide additional implementation details of the proposed methodology.
7 Additional Codebook Visualization